a) Show balance factor of every node after each insertion by creating an AVL tree with 4,2,2,0,1,4,5,9,7,1,5,3,6 number. Express all operations (single/double rotation) necessary to restore the balance. b) Create min and max heap using the same input as in part (a) by showing all the necessary steps.
Answer #1

Solution:

Explanation:

keys = 4, 2, 2, 0, -, 4, 5, 9,7, t ,S, 3,6 empty las Inserting keys into free to create Avl tree. Balamingº Faites for all tr

Insest di ② (B-2) (B=2) ② U CB-0) L ILR rotation (B3-4) Double rotation R CB=0) dout Balan factor of dis q and hide & is a so

Insest Yi (3=0) t62) (hot) B=0 (Zec (1:0) Q ų (B=0) Insert s. (B = -I) B=22) RR rotation (1=0) ( single sotation) Q (8=-1) (B

Inserta: CB-2) (B=0) Ben) Gio (B=-2) RL retation (Double @ ( 64) (B-0) (10) rofeting (Bac) sofation to Balancing factor of mo

Innasi (-)) (B-) -1) (1:1) 62 (13= 1) (B=0) Plitleding (Deulle it latior) a) 16=0) L 00) © 113--1) +B) of mode 4 is -2 so mak

CH-0) 113=0) 13--)) @ (6-0) (B:1) (2281) lu (120) (B20) (1:0) tbt (Be) Inrest 6: 2).(6=-1) (133-10 th=-1) (1=80) (B1) (3-0) (

(6) keys = 4, 2, 2, 0, 1, 4, 5, 9, 7, 1,5, 3, 6. Creating min-heap: Insert 4: Incert 2 4 the as 2<4 haapifying Lakat L and he

2 again heapitying as oca, Swap & and 2 Insest gi 2 प Swap & and Li 4 losest 4 4 u

Insors u u Insest ai 9 40 Insest z Insest di ( I and 2 swap Heapitying

Inkof si 19 2 Insest 3 G Heapifying as 3 c4 ;swap 3 and 4 2

Infert 6: 9 (finall min-heap heap) max-he ? Creating heap Inserte: ų Insert Li ha 2 Insert L 4 (& Insesto:

brises di (4 A d Iridestul lų b) Hearitying as acy, swap & and y 4 Insert si Q Heapilying as 4 C5 I swap 4 and 5 (4

again swap 4 and 5 Intestai O (2) @ 0 and a Heapifying as ocq, swap q I and Again swap 4

Again swap 5 and q. (G Insenti 8 Swap 2 and 7. 9 e Again swap 5 and 7. 9

Insest di loisest 5: 9 Swap Swap r and 5 a @ Insest 3: 9

Swap Rut 3 v 9 3 U Insert i G 5 Sway 3 and to Q

Again swap чал.4 6: Э 4 4. а (final max-heap)

I have explained each and every part with the help of images attached to it.